OpenStudy (anonymous):

Is the real number √(12.1) rational or irrational

OpenStudy (unklerhaukus):

\[√({12.1}) =\sqrt{\frac {11^2}{10}}=\frac{11}{\sqrt {10}}\]

OpenStudy (anonymous):

so wats is it

OpenStudy (unklerhaukus):

\[\frac{11}{\sqrt{10}}\]\[=\frac {11}{\sqrt{10}}\times\frac {\sqrt{10}}{\sqrt{10}}\]\[=\frac{11}{10}\sqrt {10}\] \[=\frac {11}{10}\sqrt 2\sqrt 5\]

OpenStudy (anonymous):

rational or irrational wat one of those

terenzreignz (terenzreignz):

For it to be rational, you HAVE TO BE ABLE to write it as a quotient of two integers... can you?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Therefore...? :)

OpenStudy (anonymous):

irrational

OpenStudy (anonymous):

thx

terenzreignz (terenzreignz):

no problem :)